The length of a rectangle is three inches more than the width. The area of the rectangle is 304 inches. If the
width is x, what is the length in terms of x?

Answer by checkley71(8403)

AREA=L*W OR 304=(X+3)X OR X~2+3X-304=0 OR (X-16)(X+19)=0 OR X=16 & X=-19
THUS THE LENGTH IS X+3 OR 16+3=19 IS THE LENGTH